# Pariksha : A Large-Scale Investigation of Human-LLM Evaluator Agreement on Multilingual and Multi-Cultural Data

### v1

@@ 4.4.3 Verbosity Bias
We also analyse if there is any bias by humans or LLM evaluator to pick a longer response as the better one. First, we show the distribution of the average length of winning, losing and tied responses for pairwise evaluation in Figure 9. We can see that both the winning and losing responses have similar distributions with a median length of approximately 80 words for both evaluators. Second, we also investigate the correlation between response length and the Direct Assessment scores and again find no such correlations for both the evaluators. Hence, we find no significant evidence of bias towards the length of responses against the findings of Wu and Aji (2023).
Figure 9: Response length distributions for the winning, losing and tied responses by humans and LLM evaluator in pairwise evaluations.
@@ 4.4.4 Self-Bias
Lastly, we check for self-bias by the GPT evaluator towards its own outputs across both types of evaluation. We calculate the average rank of GPT-4 in the Elo leaderboard given by both the evaluators as well as the rank of all other models which were evaluated for all the 10 languages. We find that of the 10 selected models, the average rank of GPT-4 increases by the highest amount (1.4 places) for evaluations performed by the GPT evaluator. The rationale behind using the Elo leaderboard is that simply checking the win-rate would not account for biases due to GPT evaluator giving lesser ties than humans. This indicates self-bias in GPT evaluator similar to Panickssery et al. (2024).
@@ 4.5 Human Feedback
Q1 Were the annotators able to understand the pairwise evaluation task and what problems did they face?
Majority of the annotators were able to understand the guidelines clearly and found the task simple. They also noted that linguistic acceptability played a big role in determining the easiness of the task, languages like Odia and Tamil had many grammatical errors which made it a difficult to go through responses.
Q2 Were the annotators able to understand the metrics in direct assessment?
@@
Q3 Which type of evaluation did the annotators find easier?
@@ 5 Discussion
Multilingual Performance
@@
Human-LLM Agreement
We find that LLM evaluators agree fairly well with the humans on the pairwise evaluation task in comparison to the direct assessment task. The LLM evaluator has low agreement with humans on Marathi, Bengali and Punjabi in the pairwise task and very low agreement for all languages particularly Bengali and Odia in direct assessment task. We also get feedback from the native human annotators and find direct assessment to be a harder task. On manually going through some examples, we find that humans tend to prefer outputs which are more friendly in nature, i.e., have good formatting, use colloquial language and explain with the help of examples.
We find that LLM-evaluators agree less with humans on evaluating responses with cultural nuances, suggesting that they do no possess enough cultural context to do these kinds of evaluations well. However, LLM evaluators are still able to capture general trends at a higher level as seen from the $\tau$ scores. This suggests that a human-in-the-loop or hybrid evaluation system is necessary for performing multilingual, multi-cultural evaluation.
Biases in Judgement
We look for position bias by looking at evaluator behaviour on option flipping and find no such biases. LLM evaluators are not able to detect hallucinations and pick a response even when both are hallucinated in 87% cases compared to 53% by humans. They are also found to be over-optimistic in nature. This leads to LLM evaluators having higher scores in the direct assessment task as well as fewer ties (more decisive) in pairwise evaluations task. We also look for correlations between response length and a winning response and again find no such bias. Lastly, we check for any self-bias in the GPT evaluator and find evidence of it preferring its own output.
@@ Appendix B Model Details
We list the details of all the models evaluated by us in our study in Table 2 and Table 3. We conduct evaluations on 20 indic and 10 multilingual models.
@@ Appendix C Prompt-Response Generation
All evaluated models are prompted with a system instruction followed by the query with no few-shot examples. The prompt template for each open-source model is taken from their HuggingFace model card wherever applicable, else the default Llama2-prompt is used. We instruct the models to limit their responses to 300 words and truncate the responses when necessary to make human evaluation easier, as Karya workers perform the evaluation tasks on a smartphone.
@@ D.2 Direct Assessment
‘‘name": ‘‘task quality",
@@ D.3 RTP-LX Safety Evaluation
We also do a safety evaluation study for the Hindi models using the RTP-LX de Wynter et al. (2024) Hindi dataset. Only LLM evaluators were used for this study. We used the same instruction prompt as used in direct assessment above (Figure 11 and calculate the problematic content score in the model output generations. The problematic content rubric can be seen in Figure 15.
@@ E.1 Pairwise
Figure 17: Hindi App screenshots for pairwise human evaluations
@@ E.2 Direct Assessment
Appendix F Evaluation Statistics
We perform a total of over 90k human evaluations and 30k LLM evaluations in our study. A breakdown of these statistics can be seen in Table 4. A total of 21690 datapoints (battles) are evaluated by LLMs evaluator for pairwise evaluation whereas 2880 model query-response pairs are evaluated for 3 metrics (hallucinations, task quality and linguistic acceptability) which results to 8640 datapoints in total. Hence, a total of $21690+8640=30330$ evaluations are performed by the LLM evaluator and since we annotate each datapoint by 3 humans, a total of 90990 evaluations are done by humans. In the models column, first number within parenthesis is the number of Indic-only models and the second value is the number of multilingual models under evaluation. We evaluate 12-15 models for each language (except Hindi) and 20 models for Hindi.
Appendix G Leaderboards
@@
G.1 MLE Elo Leaderboards
We report the MLE Elo leaderboards for all the 10 languages in Tables 5, 6, 7, 8, 9, 10, 11, 12, 13 and 14.
@@ G.1 MLE Elo Leaderboards
G.2 Standard Elo Leaderboards
We report the Standard Elo leaderboards for all the 10 languages in Tables 15, 16, 17, 18, 19, 20, 21, 22, 23 and 24.
@@ G.2 Standard Elo Leaderboards
G.3 Direct Assessment Leaderboards
We report the Direct Assessment leaderboards for all the 10 languages in Tables 25, 26, 27, 28, 29, 30, 31, 32, 33 and 34.
@@ G.3 Direct Assessment Leaderboards
Appendix H Detailed Agreement Scores
H.1 Percentage Agreement
@@
On average humans agree on 70% of the samples among themselves whereas the accuracy is similar albeit slightly lower for humans and LLM evaluator. We see both evaluators agree more on the non-cultural subset of queries and follow a similar trend to the Fleiss Kappa correlations reported in Table 1. A language-wise breakdown of PA scores can be seen in Figure 20. We note that humans and LLM evaluator tend to agree less on Marathi, Punjabi and Bengali.
@@
For this task, we find slightly higher agreement between humans in comparison to the pairwise evaluation and it is similar for both the query_types. However we see a decline between human and LLM evaluator agreement and it is the worse for culturally-nuanced set of queries. From Figure 20, we again find the lowest agreement on Odia and Bengali for humans and LLM evaluator, similar to the Fleiss Kappa scores.
@@
H.2 Kendall Tau
We also report a detailed language-wise breakdown of the Kendall Tau rank correlations for both evaluation tasks in Table 36. We find the correlation between human and LLM leaderboards to be higher for the pairwise evaluation. Gujarati has the highest correlation while Bengali is the lowest. For the direct assessment the scores drops and it is again the lowest for Bengali. We conclude that evaluators are better at capturing general trends in a pairwise comparison evaluation over a direct assessment task.

### v2

@@ 4.4.3 Verbosity Bias
We also analyse if there is any bias by humans or LLM evaluator to pick a longer response as the better one. From Figure 9, we observe that both humans and LLMs exhibit a slight bias towards a longer response which increases with the size difference between the responses. For LLMs, the bias remains relatively stable when the difference is 40-100 words, hovering around 0.6. Additionally, we observe that humans are slightly more biased than LLMs. We also note that the bias decreases when the difference becomes too large (over 100 words). We limit our responses to 300 words and hypothesize that such large responses might contain gibberish since LLMs are not very adept in multilingual settings. Hence, we find evidences of bias towards the length of responses by both the evaluators Wu and Aji (2023).
Figure 9: Figure showing the win fraction of a longer answer over a shorter answer (0.5 indicates random) against the difference in length of the responses in pairwise comparisons.
@@ 4.4.4 Self-Bias
Lastly, we check for self-bias by the GPT evaluator towards its own outputs across both types of evaluation. We calculate the average rank of GPT-4 in the Elo leaderboard given by both the evaluators as well as the rank of all other models which were evaluated for atleast 8-10 languages. We find that of the 11 selected models, the average rank of GPT-4 increases by the highest amount (1.4 places) for evaluations performed by the GPT evaluator (Appendix H). The rationale behind using the Elo leaderboard is that simply checking the win-rate would not account for biases due to GPT evaluator giving lesser ties than humans. This indicates self-bias in GPT evaluator similar to Panickssery et al. (2024).
@@ 4.5 Human Feedback
Q1. Were the annotators able to understand the pairwise evaluation task and what problems did they face?
Majority of the annotators were able to understand the guidelines clearly and found the task simple. They also noted that linguistic acceptability played a big role in determining the easiness of the task, languages like Odia and Tamil had many grammatical errors which made it difficult to go through responses.
Q2. Were the annotators able to understand the metrics in direct assessment?
@@
Q3. Which type of evaluation did the annotators find easier?
@@ 5 Discussion
RQ1. Are Indic LLMs able to compete with Proprietary models in respective languages?
@@
RQ2. Can LLM evaluators be used as a substitute for human evaluators in the multilingual setting? Which task format shows higher promise?
We find that LLM evaluators agree fairly well with humans on the pairwise evaluation task in comparison to the direct assessment task. The LLM evaluator has low agreement with humans on Marathi, Bengali and Punjabi in the pairwise task and very low agreement for all languages particularly Bengali and Odia in direct assessment task. We also get feedback from the native human annotators and find direct assessment to be a harder task. On manually going through some examples, we find that humans tend to prefer outputs which are more friendly in nature, i.e., have good formatting, use colloquial language and explain with the help of examples.
We find that LLM-evaluators agree less with humans on evaluating responses with cultural-nuances, suggesting that they do no possess enough cultural context to do these kinds of evaluations well. However, LLM evaluators are still able to capture general trends at a higher level as seen from the $\tau$ scores. This suggests that a human-in-the-loop or hybrid evaluation system is necessary for performing multilingual, multi-cultural evaluation.
RQ3. What are the biases that affects the evaluators’ judgements?
We look for position bias by looking at evaluator behaviour on option flipping and find no such biases. LLM evaluators are not able to detect hallucinations and pick a response even when both are hallucinated in 87% cases compared to 53% by humans. They are also found to be over-optimistic in nature. This leads to LLM evaluators having higher scores in the direct assessment task as well as fewer ties (more decisive) in pairwise evaluations task. We also look for correlations between response length and a winning response and find slight bias towards a longer response by both evaluators. Lastly, we check for any self-bias in the GPT evaluator and find evidence of it preferring its own output.
@@ Appendix B Model Details
We list the details of all the models evaluated by us in our study in Table 5 and Table 6. We conduct evaluations on 20 indic and 10 multilingual models. The models are classified based on the following criteria,
  1. 1.
    Indic: Model fine-tuned specifically for only Indian language(s).
  2. 2.
    Open-Source: Openly available models not fine-tuned specifically for Indian languages.
  3. 3.
    Proprietary: Close source models with API access and no information about the model/training data.
@@ Appendix C Prompt-Response Generation
All evaluated models are prompted with a system instruction followed by the query with no few-shot examples. The prompt template for each open-source model is taken from their HuggingFace model card wherever applicable, else the default Llama2-prompt i.e., [INST]<SYS>, is used. We instruct the models to limit their responses to 300 words and truncate the responses when necessary to make human evaluation easier, as Karya  workers perform the evaluation tasks on a smartphone.
@@ D.2 Direct Assessment
‘‘name": ‘‘task_quality",
@@ D.3 RTP-LX Safety Evaluation
We also conduct a preliminary safety evaluation study for the Hindi models using the RTP-LX de Wynter et al. (2024) Hindi dataset. Only LLM evaluators were used for this study. We used the same instruction prompt as used in direct assessment above (Figure 11) and calculate the problematic content score in the model output generations. The problematic content rubric can be seen in Figure 15.
@@ E.1 Pairwise
Figure 17: Hindi App screenshots for pairwise human evaluations.
@@ E.2 Direct Assessment
Appendix F Leaderboards
@@
F.1 MLE Elo Leaderboards
We report the MLE Elo leaderboards for all the 10 languages in Tables 7, 8, 9, 10, 11, 12, 13, 14, 15 and 16.
@@ F.1 MLE Elo Leaderboards
F.2 Standard Elo Leaderboards
We report the Standard Elo leaderboards for all the 10 languages in Tables 17, 18, 19, 20, 21, 22, 23, 24, 25 and 26.
@@ F.2 Standard Elo Leaderboards
F.3 Direct Assessment Leaderboards
We report the Direct Assessment leaderboards for all the 10 languages in Tables 27, 28, 29, 30, 31, 32, 33, 34, 35 and 36.
@@ F.3 Direct Assessment Leaderboards
Appendix G Percentage Agreement
@@
On average humans agree on 70% of the samples among themselves whereas the accuracy is similar albeit slightly lower for human-average and LLM evaluator. We see both evaluators agree more on the non-cultural subset of queries and follow a similar trend to the Fleiss Kappa correlations reported in Table 3. A language-wise breakdown of PA scores can be seen in Figure 20. We note that humans and LLM evaluator tend to agree less on Marathi, Punjabi and Bengali.
@@
For this task, we find slightly higher agreement between humans in comparison to the pairwise evaluation and it is similar for both the prompt types. However we see a decline between human and LLM evaluator agreement and it is the worse for culturally-nuanced set of queries. From Figure 20, we again find the lowest agreement on Odia and Bengali for humans and LLM evaluator, similar to the Fleiss Kappa scores.
@@
Appendix H Self-Bias
In this section we present detailed results of the self-bias analysis. Table 38 shows the average change in Elo leaderboard ranks across languages when evaluated by GPT-4 evaluator in comparison to humans. We select the 11 models which are evaluated on atleast 8-10 out of the total 10 languages. We see GPT variants rank increase the most. Gemma also shows an increase while the Aryabhatta variants and Llama-3 70B model have a big drop.
S. No. Model $\Delta$ Rank
1 GPT-4 +1.4
2 Gemma 7B +1.3
3 GPT-4o +0.6
4 GPT-35-Turbo +0.4
5 Llama-3 8B +0.1
5 Llama-2 7B +0.1
7 Mistral 7B -0.4
8 Navarasa -0.5
9 Arybhatta-GemmaOrca -1.3
10 Llama-3 70B -1.6
11 Arybhatta-GemmaUltra -1.9
Table 38: Average change in Elo Rank ($\Delta$) across languages when evaluated by GPT-evaluator in comparison to humans.